# AI Competitions and Benchmarks: How to ensure a long-lasting impact of a challenge with post-challenge paper, benchmarks and other dissemination action

### 10 Dec 2023 v1

@@ Abstract
inline,color=brown!40]Consider Shortening the title. Here is a suggestion: Towards impactful challenges: post-challenge papers, benchmarks and other dissemination actions – Gustavo inline,color=brown!40]Consider improving the English. For example, what does it mean: "The different public of…". Do you mean "The target audience of different post-challenge activities…" – Gustavo
@@
inline,color=brown!40]Four types of stakeholders play specific roles in a challenge: – Gustavo Four types of people are concerned with a challenge: the organisers, the application domain experts, the AI/Machine Learning experts on techniques relevant to the problem posed, and the participants. The participants are diverse; they can be experts from the application domain or AI, students or seasoned data scientists. Input from all these people should be gathered and made available for posterity.
@@ 4.1 Purpose of post-challenge activity
To this end, post-challenge activity is necessary, especially as challenges raw and meta outputs are often numerous. This activity should be structured to assess the potential value of those different outputs and eventually best organise them to extract meaningful analysis, discussion and conclusions. Sufficient resources must be assigned upfront to this activity. inline,color=brown!40]Consider using resources instead of manpower – Gustavo.
@@
inline,color=brown!40]It would be useful to reference to papers of the nature referred to in the itemisation, and comment on why such paper could be prototypical. Otherwise this is very dry and theoretical enumeration of classes of papers… – Gustavo
@@ 4.2.2 Participants fact sheet
At the end of the competition, it is good practice for organisers to submit a form, the “fact sheet”, to the participants to (i) have details on their best submission, which is not automatically provided by the platform, and (ii) know more about the participant themselves, who they are and how they managed their participation. It is best to advertise the form already in the last few days of the competition before participants move on to other activities and to insist that all inputs are worthwhile, even from non-top performers. The form should have a good balance between closed form Multiple Choice Questions, more straightforward to analyse, and free fields to gather specific feedback. inline,color=brown!40]MCQ may not be known to the reader. Write multiple choice questions – Gustavo The form should be anonymous by default but with the possibility of indicating the platform user name or leaving an email to continue the dialogue. The goal is to provide an overview of all participants (at least the ones who answer), while the bulk of the post-challenge activity will (and rightly so) focus on the best or most original contributions.
@@
inline,color=brown!40]are there examples that you can cite of fact sheets from competitions in which you learned something non-trivial, or how the answers have evolved in time? Otherwise, this looks like a good recipe, but with no meat to digest. – Gustavo
@@ 4.2.3 Code
inline,color=brown!40]I would emphasise that the submission of code is one of the most important components for the legacy of a challenge. It is also important to request that the code has an open-source license so that others using it don’t risk copyright infringement if used. – Gustavo
@@ 4.2.4 Competition log
inline,color=brown!40]…Reports or the discovery of cheating,… – Gustavo

### 12 Dec 2023 v2

# AI Competitions and Benchmarks: towards impactful challenges with post-challenge papers, benchmarks and other dissemination actions

@@ Abstract
Four types of stakeholders play specific roles in a challenge: the organisers, the application domain experts, the AI/Machine Learning experts on techniques relevant to the problem posed, and the participants. The participants are diverse; they can be experts from the application domain or AI, students or seasoned data scientists. Input from all these people should be gathered and made available for posterity.
@@ 4.1 Purpose of post-challenge activity
To this end, post-challenge activity is necessary, especially as challenges raw and meta outputs are often numerous. This activity should be structured to assess the potential value of those different outputs and eventually best organise them to extract meaningful analysis, discussion and conclusions. Sufficient resources must be assigned upfront to this activity.
@@
inline,color=brown!40]It would be useful to reference to papers of the nature referred to in the itemisation, and comment on why such paper could be prototypical. Otherwise this is very dry and theoretical enumeration of classes of papers… – Gustavo inline,color=cyan!40]In principle yes… – David
@@ 4.2.2 Participants fact sheet
At the end of the competition, it is good practice for organisers to submit a form, the “fact sheet”, to the participants to (i) have details on their best submission, which is not automatically provided by the platform, and (ii) know more about the participant themselves, who they are and how they managed their participation. It is best to advertise the form already in the last few days of the competition before participants move on to other activities and to insist that all inputs are worthwhile, even from non-top performers. The form should have a good balance between closed form Multiple Choice Questions, more straightforward to analyse, and free fields to gather specific feedback. The form should be anonymous by default but with the possibility of indicating the platform user name or leaving an email to continue the dialogue. The goal is to provide an overview of all participants (at least the ones who answer), while the bulk of the post-challenge activity will (and rightly so) focus on the best or most original contributions.
@@
inline,color=brown!40]are there examples that you can cite of fact sheets from competitions in which you learned something non-trivial, or how the answers have evolved in time? Otherwise, this looks like a good recipe, but with no meat to digest. – Gustavo inline,color=cyan!40]in principle yes… – David